[16] What percent of 50 is 15?

[17] In a certain diving competition, 5 judges score each dive on a scale from 1 to 10. The point value of the dive is obtained by dropping the highest score and the lowest score and multiplying the sum of the remaining scores by the degree of difficulty. If a dive with a degree of difficulty of 3.2 received scores of 7.5, 8.0, 9.0, 6.0, and 8.5, what was the point value of the dive?


[18] If 2x = 3y = 10, then 12xy =

Answer by edjones(7311)

16) 15/50=3/10=30%
.
17)
Drop 6.0 & 9.0
7.5+8.0+8.5=24
24*3.2=76.8
.
18)
2x=10
x=5
.
3y=10
y=10/3
.
12xy
=12*5*(10/3)
=4*5*10
=200
.
Ed